the square root of 2/7
Answer by edjones(8007)

square root of 2/7
=sqrt(2)/sqrt(7)
=(sqrt(2)/sqrt(7)) * (sqrt(7)/sqrt(7)) Rationalizing the denominator.
=sqrt(14)/7
.
Ed